# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 3.4.1 Accessible Spectroscopy
SPR-coupled spectrophotometry Smartphone-based Vitamin D (calcidiol) Serum Spiked from human subjects LC-MS Claim comparable LoD to control, spectra shifted right with decreasing concentration No quantitative vitamin D estimates or statistical analysis [178]
@@
Photonic crystal Benchtop Iron (ferritin) Serum Liquichek control sera Known ferritin concentrations and multiple ELISA tests Comparible recovery (>94%) and bias (by Bland-Altman analysis) to best-performing ELISA LOD higher than cutoff for anemia, lower end of deficiency [132]
@@
Sandwich iron-oxide nanoparticle immunoassay Desktop device Iron (STfR) Serum Liquichek control sera ELISA SD of 0.45 mcg/mL vs ELISA [133]
@@
UV-vis spectrophotometry Benchtop instrument Vitamin C (ascorbic acid) Serum Human subjects None Within phsyiologically-relevant concentrations Strong focus on the effectiveness of the immobilization technique, not vitamin C measurement [19]
@@
UV-vis spectrophotometry Benchtop instrument Vitamins K1 (phylloquinone), K2 (menaquinone), B6, D3 (cholecalciferol) Serum/plasma Human subjects Various sensing techniques (SWV, HPLC-MS/MS (ESI), SWAdSV, Thermal wave transport analysis, DP AdSV, SWASV, DPV, Electrochemical, Colorimetric aptasensor) Limits of detection of vitamins K1, K2, B6, and D3 are 0.075, 0.1, 0.12, and 0.15 ng/mL, respectively. Limits of quantification are 0.29, 0.3, 0.38, and 0.48 ng/mL for vitamins K1, K2, B6, and D3, respectively. Clinical gold-standard biomarkers are not used for all vitamins (except for K, where LoD is too high for deficiency); used bismuth nanoparticle embedded polypyrrole nanocomposite (rGO/pPy/Bi NC) as an optical sensing material [116]
3.5 Biofluid Analytic Methods
@@
A common approach to estimating nutritional status is by making predictions from pre-existing biofluid analysis or demographic data using ML. Such methods have been applied to derive micronutrient-specific insights [169, 94, 82, 131, 138, 104]. A summary of biofluid analytic methods can be found in Table 8.
3.5.1 Single Micronutrient Malnutrition Detection
@@
Some studies focus on detecting deficiency of a specific micronutrient. Two such papers target iron status [105, 138], while a third targets vitamin D [131]. Luo et al. [105] used hospital outpatient data collected over three months to predict whether a patient had normal or abnormal ferritin (iron) status. The collected data included age, sex, ferritin test results (used as targets), and other ‘predictor’ tests that were conducted in the main hospital lab only (n=5128). These authors considered the broader clinical usefulness of ML-powered insights, claiming that “predicted ferritin results may sometimes better reflect underlying iron status than measured ferritin” [105]. This conclusion was based on dual independent review by pathologists on 26 selected cases where measured and predicted ferritin were ‘highly discrepant’. They propose that predicted levels could be used to flag lab-measured ferritin for further review. Pullakhandam and McRoy [138] use gradient boost on NHANES complete blood count (CBC) data to classify and explain IDA (n=19995). They found that the most critical features for IDA are low levels of hemoglobin, higher age, and a higher red blood cell distribution width.
@@
For the prediction of vitamin D deficiency, Patino-Alonso et al. [131] applied ML to anthropomorphic data of older Europeans (35-75 y/o; 50/50 males/females) given anthropomorphic features. 501 participants contributed their “waist circumference (WC), body mass index (BMI), waist-to-height ratio (WHtR), body roundness index (BRI), visceral adiposity index (VAI), and the Clinical University of Navarra body adiposity estimator (CUN-BAE) for body fat percentage”. Vitamin D as 25(OH)D was measured by immunoassay and the threshold of deficiency was set to be 20 ng/mL (34.7% prevalence). We note that this threshold more closely aligns with insufficiency (Table 18). Logistic regression analysis found that the most significant features differed by sex. All but CUN-BAE were associated with vitamin D deficiency in males, while only CUN-BAE was associated in females. ML models for deficiency prediction were trained on each feature individually. The authors discovered that Naive Bayes was the top performer by AUC for WC, BMI, WHtR, and BRI but was bested by logistic regression for VAI and CUN-BAE.
3.5.2 Multiple Micronutrient Detection
@@
Since it is rare for micronutrient imbalance to occur in isolation, researchers have studied the ability to predict malnutrition of multiple micronutrients [169, 94]. Truijen et al. [169] focus on micronutrient malnutrition in older populations, citing how malnutrition in older adults is often diagnosed too late despite the existence of screening methods. The goal of the study was to classify each sample as having either no micronutrient deficiency or one or more deficiencies among vitamins C, B6, B12, selenium, and zinc, confirmed by blood tests. These particular micronutrients were selected because they interact less with each other (we add that B-vitamins do interact; Table LABEL:tab:char-vitamins-water-soluble), were among the most prevalent deficiencies, had clinically relevant cutoff points for deficiency. Logistic regression was applied to routine biochemical and diagnostic data from 9 years of United Kingdom NDNS for ages $\geq$50 (n=1518). This dataset suffered from ethnic disparities, with the authors noting limited generalizability to ethnic groups due to $\geq$95of NDNS participants being white. Kurstjens et al. [94] aimed to develop an ML algorithm to assess risk of low body iron storage (ferritin plasma levels) in anemic primary care patients using CBC and CRP tests. Two algorithms were developed, each based on laboratory ferritin results from different chemistry analyzers (from Siemens and Roche). Interestingly, the two most important features were both derived from CBC test results (Table 8). The authors took an important step to consider how such a model could assist a clinician by asking professionals to indicate whether a patient had low ferritin based on CBC and CRP, with and without algorithm results. The found that the algorithm alone was more accurate than both scenarios. Detection of low vitamin B12 and B9 levels were also considered, but this yielded poor results with AUCs of 0.52 and 0.57 respectively.
3.5.3 Adjusting Biomarkers for Inflammation
@@
When conducting biochemical analysis for micronutrient assessment, a common issue is the impact of inflammation on biomarker measurement. One R package aims to solve this problem and improve interpretability by adjusting biomarkers of micronutrients in the context of inflammation [104]. The package implements inflammation adjustment for retinol-binding protein, serum retinol, serum ferritin, sTfR, and serum zinc, using acid glycoprotein (AGP) and/or CRP as biomarkers for inflammation. The authors have also published a paper describing a procedure on when and how to apply their technique [104].
Table 8: Analytic Methods
@@
Logistic regression Normal or abnormal ferritin Hospital outpatient data "total iron-binding capacity, mean cell hemoglobin, and mean cell hemoglobin concentration" [105] Ferritin test results AUC of 0.97 Predictions could be used to flag lab ferritin for review [105]
@@
Gradient boost IDA US NHANES (CBC and serum ferritin) dataset and Kenyan nutrition dataset Low blood level of hemoglobin, higher age, and higher red blood cell distribution width Serum ferritin PR AUC of 0.87 and recall/sensitivity of 0.98 and 0.89 for original and test dataset, respectively Heavy class imbalance (4.9% IDA vs. 95.1% non-IDA) [138]
@@
Logistic regression, naive bayes, and random forest Vitamin D deficiency Anthropormorphic measurements of older Europeans CUN-BAE for females, all others for males Blood 25(OH)D by immunoassay Max AUC of $\sim$0.53 for all features; LR best for VAI and CUN-BAE, NB for all others Did not assess predictive ability of multiple features at once [131]
@@
Logistic regression Presence of micronutrient deficiency UK NDNS, ages $\geq$50 Low protein, energy intake, TC, hemoglobin, HbA1c, ferritin, vitamin D and high CRP Blood test results for Vitamins C, B6, B12, selenium, and zinc AUC of 0.79 $\geq$95of NDNS participants were white [169]
@@
Random forest Classify low body iron storage (plasma ferritin) CBC and CRP tests in anemic primary care patients Mean corpuscular hemoglobin and mean corpuscular volume Ferritin results from two laboratory chemistry analyzers (two separate models) AUC of 0.9 and 0.92 for each model, models were more accurate than professionals with and without access to results Attempted Vitamin B12 and B9 deficiency detection with poor results (AUCs of 0.52 and 0.57) [94]
3.6 Biofluid Analysis Limitations
The largest general limitation of existing biofluid-based assessments is that the most effective require an invasively-collected biosample. Methods that indicate in-body status of a micronutrient often utilize a blood sample for analysis. Clinical research on micronutrients suggests that sweat and saliva may not accurately reflect micronutrient status compared to blood, yet many methods continue to overlook this discrepancy. We also observe that biofluid-based assessments are also more specialized in nature. Frequently, a bespoke assay, device, or sensor is implemented for the assessment of only a single micronutrient. This is understandable considering the inherent difficulties in detecting and quantifying micronutrients that have unique metabolic pathways, are present in such limited quantities, and play different roles in bodily function. We find that current methods do not consistently test the clinical biomarker for a given micronutrient (e.g. measuring Vitamin D via calciol instead of the gold-standard calcidiol/25(OH)D). Even when the proper biomarker is targeted, the device itself may be evaluated on concentration intervals that are not pertinent to the spectrum of deficiency to excess (e.g. the linear range of a sensor may be able to indicate sufficient status, but the LoD is too high to infer deficiency). Some research fails to include a clinically-relevant biochemical test (e.g. ELISA or HPLC) for their target micronutrient as a scientific control. Such oversights actively limit the clinical value, and therefore the real-world utility, of the proposed solution. Therefore, we have taken note of when this is the case, both in the main body of the text and the above tables.
@@ 3.6 Biofluid Analysis Limitations
Moving to the specific approaches, we find that clinical biochemical analysis is invasive, expensive to analyze, and the methods of analysis and thresholds for imbalance are debated [142]. Additionally, biomarkers are generally sensitive but not specific, and their analysis requires considering an extensive list of factors that alter the ability of a marker to indicate nutrient status (such as inflammation, disease, or medications) [43].
@@
Although there are several wearable, smartphone, and point of care assay devices, their accessibility in some cases is limited by specialized assay chips and most require invasive biofluids for analysis. Saliva and sweat have been proposed as alternatives to blood, but their ability to reflect in-body status of a given micronutrient is often unknown, debated, or disproved (Section 2.2). One survey offers additional insight into the issues faced by wearable sweat sensors: low sweat rates, sample evaporation, skin contamination impacting sweat content, and the difficulty to access fresh sweat [15]. Assays and electrochemical devices also suffer from being more specialized in nature (complicating manufacture and integration) and not utilizing more ubiquitous methods of health monitoring such as smartphones and smartwatches. Most work is limited to the measurement of one vitamin, claiming that simultaneous detection of multiple vitamins, especially both water and fat-soluble vitamins at once, is more limited. There is also a common need for a buffer or other solution to serve as an electrolyte for the electrochemical analysis of a vitamin. Like reagents in an assay, this adds complexity to the measurement process and creates a barrier to widespread adoption.
@@
The field of spectroscopy shows great potential (especially IR and Raman), though we argue that there is not yet enough accessible, micronutrient-specific research that provides insights into in-body status. Most work in this area required benchtop analyzers instead of on-body approaches. With the latter however, one must take great care not to harm a user with the spectroscopic approach. UV spectroscopy, for example, requires the application of UV light and could cause skin damage.

### v2

@@ 3.4.1. Accessible Spectroscopy
SPR-coupled spectrophotometry Smartphone-based Vitamin D (calcidiol) Serum Spiked from human subjects LC-MS Claim comparable LoD to control, spectra shifted right with decreasing concentration No quantitative vitamin D estimates or statistical analysis (Walter et al., 2020)
@@
Photonic crystal Benchtop Iron (ferritin) Serum Liquichek control sera Known ferritin concentrations and multiple ELISA tests Comparible recovery (¿94%) and bias (by Bland-Altman analysis) to best-performing ELISA LOD higher than cutoff for anemia, lower end of deficiency (Peterson et al., 2014)
@@
Sandwich iron-oxide nanoparticle immunoassay Portable Iron (STfR) Serum Liquichek control sera ELISA SD of 0.45 mcg/mL vs ELISA (Peterson et al., 2015)
@@
UV-vis spectrophotometry Benchtop Vitamin C (ascorbic acid) Serum Human subjects None Within phsyiologically-relevant concentrations Strong focus on the effectiveness of the immobilization technique, not vitamin C measurement (Bi et al., 2016)
@@
UV-vis spectrophotometry Benchtop Vitamins K1 (phylloquinone), K2 (menaquinone), B6, D3 (cholecalciferol) Serum/plasma Human subjects Various sensing techniques (SWV, HPLC-MS/MS (ESI), SWAdSV, Thermal wave transport analysis, DP AdSV, SWASV, DPV, Electrochemical, Colorimetric aptasensor) Limits of detection of vitamins K1, K2, B6, and D3 are 0.075, 0.1, 0.12, and 0.15 ng/mL, respectively. Limits of quantification are 0.29, 0.3, 0.38, and 0.48 ng/mL for vitamins K1, K2, B6, and D3, respectively. Clinical gold-standard biomarkers are not used for all vitamins (except for K, where LoD is too high for deficiency); used bismuth nanoparticle embedded polypyrrole nanocomposite (rGO/pPy/Bi NC) as an optical sensing material (Mughal et al., 2024)
3.4.2. Comparison
A wide range of spectroscopic methods have been applied for the quantification of different micronutrients, making it difficult to compare each approach. Iron assessment appears to be feasible in serum, but further development would be useful to increase the accessibility of a spectroscopic approach that targets this nutrient. Furthermore, clinical research could consider alternative biosamples for iron, to increase the noninvasiveness of spectroscopic techniques. Studies using spectrophotometric methods have yet to be thoroughly validated so as to indicate the utility of spectrophotometry over any alternative method.
3.5. Biofluid Analytic Methods
@@
AI and ML can be used to detect micronutrient levels in individuals. A common approach to estimating nutritional status is by making predictions from pre-existing biofluid analysis or demographic data using ML. Such methods have been applied to derive micronutrient-specific insights (Truijen et al., 2021; Kurstjens et al., 2022; Khan et al., 2021; Patino-Alonso et al., 2022; Pullakhandam and McRoy, 2024; Luo et al., 2023). A summary of biofluid analytic methods can be found in Table 9, where we observe a common trend of utilizing classical machine learning algorithms (logistic regression, gradient boost, naive bayes, random forest) for predictive modeling.
3.5.1. Single Micronutrient Malnutrition Detection
@@
Some studies focus on detecting deficiency of a specific micronutrient. Two such papers investigate iron status (Luo et al., 2016; Pullakhandam and McRoy, 2024), while a third examines vitamin D (Patino-Alonso et al., 2022). Luo et al. (2016) used hospital outpatient data collected over three months to predict whether a patient had normal or abnormal ferritin (iron) status using logistic regression. The collected data included age, sex, ferritin test results (used as markers), and other ‘predictor’ tests that were conducted in the main hospital lab only (n=5128, sex breakdown not reported). These authors considered the broader clinical usefulness of ML-powered insights, claiming that “predicted ferritin results may sometimes better reflect underlying iron status than measured ferritin” (Luo et al., 2016). This conclusion was based on dual independent review by 2 pathologists on 26 selected cases where measured and predicted ferritin were ‘highly discrepant’. They propose that predicted levels could be used to flag lab-measured ferritin for further review. Pullakhandam and McRoy (2024) use gradient boost on NHANES complete blood count (CBC) data to classify and explain IDA (n=19995, sex breakdown not reported). They found that the most critical features for IDA are low levels of hemoglobin, higher age, and a higher red blood cell distribution width.
@@
For the prediction of vitamin D deficiency, Patino-Alonso et al. (2022) applied ML (logistic regression, naive bayes, and random forest) to anthropomorphic data of older Europeans (35-75 y/o; 50/50 males/females) given anthropomorphic features. A total of 501 participants contributed their “waist circumference (WC), body mass index (BMI), waist-to-height ratio (WHtR), body roundness index (BRI), visceral adiposity index (VAI), and the Clinical University of Navarra body adiposity estimator (CUN-BAE) for body fat percentage”. Vitamin D as 25(OH)D was measured by immunoassay and the threshold of deficiency was set to be 20 ng/mL (34.7% prevalence). We note that this threshold more closely aligns with insufficiency (Table LABEL:tab:_assessment). Logistic regression analysis found that the most significant features differed by sex. All but CUN-BAE were associated with vitamin D deficiency in males, while only CUN-BAE was associated in females. ML models for deficiency prediction were trained on each feature individually. The authors discovered that Naive Bayes was the top performer by AUC for WC, BMI, WHtR, and BRI but was bested by logistic regression for VAI and CUN-BAE.
3.5.2. Multiple Micronutrient Detection
@@
Since it is rare for micronutrient imbalance to occur in isolation, researchers have studied the ability to predict malnutrition of multiple micronutrients (Truijen et al., 2021; Kurstjens et al., 2022). Truijen et al. (2021) focus on micronutrient malnutrition in older populations, citing how malnutrition in older adults is often diagnosed too late despite the existence of screening methods. The goal of the study was to use logistic regression to classify each sample as having either no micronutrient deficiency or one or more deficiencies among vitamins C, B6, B12, selenium, and zinc, confirmed by blood tests. These particular micronutrients were selected because they interact less with each other (we add that B-vitamins do interact; Table LABEL:tab:char-vitamins-water-soluble), were among the most prevalent deficiencies, had clinically relevant cutoff points for deficiency. Logistic regression was applied to routine biochemical and diagnostic data from 9 years of United Kingdom NDNS for ages $\geq$50 (n=1518, 57.2% female). This dataset suffered from ethnic disparities, with the authors noting that $\geq$95of NDNS participants were white.
Kurstjens et al. (2022) aimed to develop a random forest algorithm to assess risk of low body iron storage (ferritin plasma levels) in anemic primary care patients using CBC and CRP test results from 3 medical laboratories (n=2,935, $\sim$1,493 female. Two algorithms were developed, each based on laboratory ferritin results from different chemistry analyzers (from Siemens and Roche). Interestingly, the two most important features were both derived from CBC test results (Table 9). The authors took an important step to consider how such a model could assist a clinician by asking 4 professionals to indicate whether a patient had low ferritin based on CBC and CRP, with and without algorithm results. The found that the algorithm alone was more accurate than both scenarios. Detection of low vitamin B12 and B9 levels were also considered, but this yielded poor results with AUCs of 0.52 and 0.57 respectively.
3.5.3. Adjusting Biomarkers for Inflammation
@@
When conducting biochemical analysis for micronutrient assessment, a common issue is the impact of inflammation on biomarker measurement. One R package aims to solve this problem and improve interpretability by adjusting biomarkers of micronutrients in the context of inflammation (Luo et al., 2023). The package implements inflammation adjustment for retinol-binding protein, serum retinol, serum ferritin, sTfR, and serum zinc, using acid glycoprotein (AGP) and/or CRP as biomarkers for inflammation. The authors have also published a paper describing a procedure on when and how to apply their technique (Luo et al., 2023).
Table 9. Analytic Methods
@@
Logistic regression Normal or abnormal ferritin Hospital outpatient data ”total iron-binding capacity, mean cell hemoglobin, and mean cell hemoglobin concentration” from Luo et al. (2016) Ferritin test results AUC of 0.97 Predictions could be used to flag lab ferritin for review (Luo et al., 2016)
@@
Gradient boost IDA US NHANES (CBC and serum ferritin) dataset and Kenyan nutrition dataset (for evaluation) Low blood level of hemoglobin, higher age, and higher red blood cell distribution width Serum ferritin Precision AUC of 0.87 (training); recall of 0.98 (training) and 0.89 (evaluation) Heavy class imbalance (4.9% IDA vs. 95.1% non-IDA) (Pullakhandam and McRoy, 2024)
@@
Logistic regression, naive bayes, and random forest Vitamin D deficiency Anthropormorphic measurements of older Europeans CUN-BAE for females, all others for males Blood 25(OH)D by immunoassay Max AUC of $\sim$0.53 for all features; LR best for VAI and CUN-BAE, NB for all others Did not assess predictive ability of multiple features at once (Patino-Alonso et al., 2022)
@@
Logistic regression Presence of micronutrient deficiency UK NDNS, ages $\geq$50 Low protein, energy intake, TC, hemoglobin, HbA1c, ferritin, vitamin D and high CRP Blood test results for Vitamins C, B6, B12, selenium, and zinc AUC of 0.79 $\geq$95of NDNS participants were white (Truijen et al., 2021)
@@
Random forest Classify low body iron storage (plasma ferritin) CBC and CRP tests in anemic primary care patients Mean corpuscular hemoglobin and mean corpuscular volume Ferritin results from two laboratory chemistry analyzers (two separate models) AUC of 0.9 and 0.92 for each model, models were more accurate than professionals with and without access to results Attempted Vitamin B12 and B9 deficiency detection with poor results (AUCs of 0.52 and 0.57) (Kurstjens et al., 2022)
3.5.4. Comparison
Luo et al. (2016) and Kurstjens et al. (2022) both utilize data obtained from patients during the course of their health care, whereas other sources of data came from surveys (e.g. the US NHANES). Collecting data during health care can result in a large amount of data to mine for micronutritonal insights without needing to rely on national surveys (which may not contain a wide variety of features) or conducting an independent assessment of a population. The prediction of iron (as ferritin) appears to be more successful than the prediction of other biomarkers. This is likely due to iron’s large influence on hemoglobin, which was an important feature for all ML models that were trained to predict ferritin status (Luo et al., 2016; Kurstjens et al., 2022; Pullakhandam and McRoy, 2024).
3.6. Biofluid Analysis Limitations
The largest general limitation of existing biofluid-based assessments is that methods effectively indicating in-body status of a micronutrient often utilize an invasively-collected biosample (i.e. blood) for analysis. While noninvasive biosamples such as sweat and saliva were applied, the clinical research on micronutrients suggests that these samples may not accurately reflect micronutrient status compared to blood. As such, validation studies are required. We also observe that biofluid-based assessments are more specialized in nature. Frequently, a bespoke assay, device, or sensor is implemented for the assessment of only a single micronutrient. This is understandable considering the inherent difficulties in detecting and quantifying micronutrients that have unique metabolic pathways, are present in such limited quantities, and play different roles in bodily function. Furthermore, there is also a common need for a specialized reagent, buffer, or electrolytic solution to enable the analysis of each micronutrient (especially for assays and electrochemical methods). It is for this reason that multiplexed or simultaneous detection of multiple micronutrients is limited, despite their potential to enable more holistic status monitoring. However, we acknowledge that combining several specialized assays and unique sensor designs adds complexity to the measurement process, creating barriers to commercialization as well as widespread adoption.
On the topic of commercialization, we posit that the lack of commercial, multiplexed, or PoC technologies for micronutrient assessment is a result of both the ingrained reliance on laboratory blood testing within health care institutions as well as the expensive and thorough human subjects evaluations necessary to bring an effective assessment technique to market. We find that current methods do not consistently test the clinical biomarker for a given micronutrient (e.g. measuring vitamin D via calciol instead of the optimal reference standard calcidiol a.k.a. 25(OH)D). Even when the proper biomarker is examined, the device itself may be evaluated on concentration intervals that are not pertinent to the clinical spectrum of deficiency to excess (e.g. the linear range of a sensor may be able to indicate sufficient status, but the LoD is too high to infer deficiency). Again, we recognize that some methods are specifically designed to accurately detect a deficient status in a resource-constrained environment, or assess micronutrients where excess does not pose a significant health risk. In these cases, a lack of sensitivity in the upper spectrum is less of an issue. Some research fails to include a clinically relevant biochemical test (e.g. ELISA or HPLC) for their target micronutrient as a scientific control. By contrast, a non-clinically relevant biochemical test either does not follow best practices in clinical nutrition, targets an alternative biomarker, or uses an alternative process from what is typically considered acceptable in modern clinical practice. Such a non-clinically relevant test can also simply rely upon supplementation rather than any substantive evaluation of micronutrient status (see below). Each of these oversights actively limits the clinical value, and therefore the real-world utility, of the proposed solution.
The lack of a clinically relevant biochemical test is confounded by the prevalence of supplementation-based experiments, where an assessment is conducted on a patient before, during, and/or after intake of the target micronutrient. The prescribed intake may be as specific as a supplement pill or as general as a food that is known to contain high amounts of a micronutrient. Because of individual variance in the absorption of different micronutrients, it is difficult to know the true impact of micronutrient intake, and therefore the practical accuracy of a status assessment, without also conducting a relevant biochemical test. Lastly, the study designs themselves often feature low sample sizes (1-5 participants) that make it difficult to assess the analytical validity of each method. We recognize that design-oriented research faces a larger hurdle when attempting to conduct human-subjects experiments that are clinical in nature, and several studies echo this point. Biofluid analytic methods have sample sizes that are (necessarily) much larger, but share issues related to the equitable representation of demographic groups within their sample. Similar to the impact of omitting a relevant biochemical test, we argue that an emphasis on new technologies over more rigorous and larger-scale validation studies makes it difficult to judge the true effectiveness of a novel assessment method.
@@ 3.6. Biofluid Analysis Limitations
Moving to the specific approaches, we find that clinical biochemical analysis is invasive, expensive to analyze, and the methods of analysis and thresholds for imbalance are debated (Reber et al., 2019). Additionally, biomarkers are generally sensitive but not specific, and their analysis requires considering an extensive list of factors that alter the ability of a marker to indicate nutrient status (such as inflammation, disease, or medications) (Elmadfa and Meyer, 2014).
@@
Although there are several wearable, smartphone, and point of care assay devices, their accessibility in some cases is limited by specialized assay chips and most require invasive biofluids for analysis. Saliva and sweat have been proposed as alternatives to blood, but their ability to reflect in-body status of a given micronutrient is often unknown, debated, or disproved (Section 2.2). One survey offers additional insight into the issues faced by wearable sweat sensors: low sweat rates, sample evaporation, skin contamination impacting sweat content, and the difficulty to access fresh sweat (Bariya et al., 2018). Assays and electrochemical devices also suffer from being more specialized in nature (complicating manufacture and integration) and not utilizing more ubiquitous methods of health monitoring such as smartphones and smartwatches.
@@
The field of spectroscopy shows great potential (especially IR and Raman), though we argue that there is not yet enough accessible, micronutrient-specific research that provides insights into in-body status. Most work in this area required benchtop analyzers instead of on-body approaches. With the latter however, one must take great care not to harm a user with the spectroscopic approach. UV spectroscopy in particular requires the application of UV light, which is broken into three types depending on its wavelength. Exposure to UVA (400 - 320 nm) and UVB (320 - 280 nm) radiation can cause damage to DNA and result in cancer, because of how this energy penetrates the skin (D’Orazio et al., 2013). UVC radiation from sunlight (280 - 100 nm) does not possess these same dangers, especially because it is absorbed by the skin, but direct exposure to artificial UVC radiation can still cause burns on the skin and eyes (United States Food & Drug Administration, 2020). The exact amount of UV radiation which is considered harmful varies with the area of exposure, amount of melanin content in the skin, and more, so we direct the reader to Table 2 in D’Orazio et al. (2013) for more information.